- 187: CO2, Site tower, Thu 28-Oct-1999 12:30:37 CDT, Licor zero
At ~ 1200 the Licor was zeroed for ~ 15 minutes
- 182: CO2, Site , Wed 27-Oct-1999 11:21:50 CDT, Licor zero
At 07:25 , 27 Oct''99 Licor zero cannister attached.
At 08:00 , 27 Oct''99 Licor zero cannister removed.
- 175: CO2, Site , Tue 26-Oct-1999 13:09:08 CDT, Zero Licor
Yesterday between 1515 and 1645 the zero cannister was attached to the Licor
intake.
Today between 1030 and 1115 the zero cannister was attached to the Licor intake.
The signal during operation was ~ 2.36
The signal during zero was ~ -0.214
- 170: CO2, Site tower, Sun 24-Oct-1999 13:01:17 CDT, Zeroed Licor
Between 1230 and 1300 a zero canister was attached to the intake of the
Licor CO2 sensor.
Front panel display for zero = - 0.2
Front panel display for ambient = + 2.4 to 2.5
The pump was running well but I couldn't check the flow as the flow guage
was missing.
Both cal cylinders were empty.
- 141: CO2, Site tower, Tue 19-Oct-1999 19:07:45 CDT, Mounted new h2o/co2 sensor at 20m
ABLE (Argonne) provided a replacement for the bad h2o/co2 sensor
at 20m. We mounted this on the tower this evening, raising the 20m
boom around 1710 CDT and lowering it at 1755 CDT.
At this time we removed the KH2O sensor for remounting on Station 6.
- 137: CO2, Site tower, Mon 18-Oct-1999 16:40:30 CDT, Removed 20m h2o/co2 sensor
We have given up on fixing the ANL h2o/co2 sensor on the 20m boom.
We raised the boom at 15:20 to remove it and lowered the boom at
16:05 (after adding another collar to hold the boom pivot in
position).
- 125: CO2, Site tower, Sun 17-Oct-1999 19:17:25 CDT, CO2 back on at 20 meters
The CO2 sensor was remounted at 20 meters after a preliminary
test with the new 5 volt supply. We left the krypton there
as the primary h2o sensor. The boom was lowered at 18:25 CDT.
- 120: CO2, Site tower, Sun 17-Oct-1999 14:46:38 CDT, work on CO2 sensor
It appears that the problem with the CO2/H2O sensor was
in the 5 volts supply. I modified their power supply so
that we could use a different 5 volt source. The sensor
was deployed for testing at the 20 meter height. After
we collect some data we will decide if it is good or bad.
The sensor was deployed at ~13:50 local time.
- 106: CO2, Site tower, Sat 16-Oct-1999 08:46:59 CDT, 20m co2 status
I ran some tests on the 20m co2 sensor. Based on feedback from
Dave Cook at Argone, it appears the sensor is bad. At one time,
it appeared the 5 volt supply was the problem; however, tests
with the other co2 supply showed this not to be the case. We will
need to send the unit back to Dave Cook.
- 105: CO2, Site tower, Sat 16-Oct-1999 08:43:48 CDT, 40m co2 power supply taken down
The power supply unit for the 40m co2 sensor was removed
last night to run tests with the questionable co2 sensor
deployed at 20m. No data is being lost since ragwort has been
turned off.
- 95: CO2, Site tower, Thu 14-Oct-1999 09:55:16 CDT, No licor cal gases
All of the "lo-cal" CO2 calibration gas was used after the first day that
the Licor was deployed - about 10/7? I had identified a leak in the solenoid
valve which controlled this line in Tony's box and had tried to fix it, but
apparently was unsuccessful. There still was plenty of "hi-cal" gas until
sometime during the last 3 days. Yesterday morning, the "hi-cal" cylinder
also was empty.
Both cylinders are now off. We probably should disconnect the control circuit
to let the licor sniff real air all the time, but I haven't done this yet.
I don't know how we'll generate a calibration now - perhaps we should manually
attach soda lime to get a zero from time-to-time. I've put a call in to
Tony.
- 78: CO2, Site tower, Mon 11-Oct-1999 16:33:54 CDT, 20m H2O/CO2 reinstalled
We just put the ATDD sensor back up. Dave Auble simply tightened connectors
inside the electronics/power supply box and removed two passive filters
added by ANL. The data look almost believable now.
- 77: CO2, Site tower, Mon 11-Oct-1999 13:26:40 CDT, 20m boom up/CO2 removed
We've just removed the 20m H2O/CO2 sensor so that Dave Auble can look at
it today. This was from about 12:30-13:00
- 59: CO2, Site tower, Fri 08-Oct-1999 17:07:37 CDT, co2.20m fixed itself?
co2.20m appears to have a reasonable signal now, despite no action being
taken on our part. It definitely has been offscale or spiking for the
previous several days. We'll keep watching it.
- 54: CO2, Site tower, Fri 08-Oct-1999 09:03:30 CDT, approximate licor calibration
A crude fit between the Licor and co2.40m gives:
CO2 = (Licor + 2.4)*140
in mg/m3 of CO2. To convert to ppm:
mg/m3 * g/1000mg * mole/44g * 22.4 lit/mole * m3/10^3 lit * 10^6ppm ~ 0.5
With these factors, the cal-gas values are about:
357ppm; 330ppm; 364ppm
I will enter these values in prep.config.
- 52: CO2, Site tower, Fri 08-Oct-1999 08:44:28 CDT, Licor co2 installed
Yesterday, I found that the Licor problem was wrong pins connected to the
CO2 analog output channel and corrected this problem about noon. Signal +
was connected to power ground (pin7) and Signal return to open (pin4).
The signal now appears to be working. At that time, I also opened the cal
gas cylinders and installed the valve control lines. I see 2 calibrations
in the data at about 4PM and 4AM last night.
- 48: CO2, Site tower, Tue 05-Oct-1999 19:41:05 CDT, co2 turned off
turned off Licor (includeing pump and cal gases), since it has been reporting
bad values. This done about 6pm local. It has been running bad for the last
day and a half.